# Robust MultiSpecies Agricultural Segmentation Across Devices, Seasons, and Sensors Using Hierarchical DINOv2 Models

### v2

@@ Abstract
Reliable plant species and damage segmentation for herbicide field research trials requires models that can withstand substantial real‑world variation across seasons, geographies, devices, and sensing modalities. Most deep learning approaches trained on controlled datasets fail to generalize under these domain shifts, limiting their suitability for operational phenotyping pipelines.
This study evaluates a segmentation framework that integrates vision foundation models (DINOv2) with hierarchical taxonomic inference to improve robustness across heterogeneous agricultural conditions. We train on a large, multi‑year dataset collected in Germany and Spain (2018–2020), comprising 14 plant species and 4 herbicide damage classes, and assess generalization under increasingly challenging shifts: temporal and device changes (2023), geographic transfer to the United States, and extreme sensor shift to drone imagery (2024).
Results show that the foundation‑model backbone consistently outperforms prior baselines, improving species‑level F1 from 0.52 to 0.87 on in‑distribution data and maintaining significant advantages under moderate (0.77 vs. 0.24) and extreme (0.44 vs. 0.14) shift conditions. Hierarchical inference provides an additional layer of robustness, enabling meaningful predictions even when fine‑grained species classification degrades (family F1: 0.68, class F1: 0.88 on aerial imagery). Error analysis reveals that failures under severe shift stem primarily from vegetation–soil confusion, suggesting that taxonomic distinctions remain preserved despite background and viewpoint variability.
The system is now deployed within BASF’s phenotyping workflow for herbicide research trials across multiple regions, illustrating the practical viability of combining foundation models with structured biological hierarchies for scalable, shift‑resilient agricultural monitoring.
@@ 7 Conclusions
This study examines the applicability of self-supervised vision foundation models, specifically DINOv2, for multi-species weed segmentation and damage identification under substantial domain shift. Unlike most prior research, which primarily assesses in-domain performance, the proposed approach is evaluated across five growing seasons, three countries, and multiple imaging platforms, including smartphones, handheld cameras, and drones. These experiments encompass significant variability in sensor characteristics and acquisition conditions, including transitions from ground-based to aerial imagery. The results suggest that self-supervised representations, combined with hierarchical taxonomic inference, can offer robustness under such operational variability. Evaluation on herbicide field trials involving more than 70 species further indicates that the model can accommodate temporal, geographic, and technological heterogeneity, pointing to its potential suitability for scalable automated agricultural monitoring.
The three-stage evaluation protocol yields three principal findings. First, the DINOv2-based architecture consistently outperforms a DeepLabV3+ baseline across all considered conditions. On the BASE dataset, species-level F1 increases from 0.52 to 0.87, and this advantage is maintained under moderate domain shift (REALITY: 0.77 vs. 0.24) as well as under substantial changes in sensor modality (DRONE: 0.44 vs. 0.14). Second, hierarchical inference contributes meaningfully to interpretability when fine-grained performance declines: although species-level accuracy decreases for aerial imagery, the system retains comparatively strong performance at higher taxonomic levels (family F1: 0.68; class F1: 0.88). Third, errors under extreme domain shift primarily involve confusion between vegetation and soil rather than inter-species misclassification, indicating that species-level distinctions remain preserved and that challenges arise chiefly from background separation.
The incorporation of the system into BASF’s phenotyping pipeline provides initial evidence of its applicability to real-world crop and weed monitoring across diverse geographic settings. While the present study emphasizes evaluating foundation models under substantial temporal, geographic, and sensor variability—including differences in years (2018–2024), countries (Germany, Spain, United States), and imaging devices—operational deployments are expected to benefit from incorporating training data spanning all relevant regions and sensors.
Several limitations highlight directions for future research. The geographic distribution of training data is restricted to Germany and Spain (2018–2020), likely contributing to lower performance at U.S. sites, although the model still exceeds baseline results. Damage classification remains difficult for early-stage symptoms (initial damage F1: 0.28) and for previously unseen damage types (e.g., leaf-curling in weeds, F1: 0.05), reflecting both annotation ambiguity and limited coverage of damage patterns. Performance degradation in drone imagery appears to stem primarily from vegetation–soil separation challenges rather than species-level confusion, though annotation inconsistencies may contribute. Incorporating aerial imagery into the training corpus may therefore strengthen fine-grained classification from drones. Furthermore, the absence of explicit uncertainty quantification limits the system’s capacity for automated quality control, including automatically flagging predictions requiring human verification.
Future work should prioritize three main areas: (1) broadening training data diversity to include wider geographic regions, additional sensor modalities—particularly aerial platforms—and a larger variety of damage patterns with refined annotation procedures; (2) developing uncertainty estimation methods to support automated quality control and confidence-aware decision-making; and (3) conducting extended temporal validation across multiple growing seasons to assess long-term robustness. These enhancements would further improve both the stability and operational suitability of the system while retaining the advantages associated with self-supervised foundation models.
Overall, this work demonstrates that combining self-supervised pretraining with structured domain knowledge can yield agricultural monitoring systems capable of generalizing across meaningful sources of variation encountered under realistic field conditions. By systematically evaluating temporal, geographic, device-related, and sensor-level domain shifts—rather than limiting analysis to in-domain performance—the study provides evidence that such approaches may satisfy the requirements of real-world herbicide trial environments. The methodological principles illustrated here, including the integration of hierarchical taxonomic inference and rigorous cross-domain evaluation, offer a foundation for developing reliable and scalable precision agriculture systems.

### v4

# Mitigating Domain Drift in Multi Species Segmentation with DINOv2: A Cross-Domain Evaluation in Herbicide Research Trials

@@ Abstract
We evaluate the effect of domain drift on plant species and herbicide‑damage segmentation in realistic, complex agricultural scenarios, and quantify how vision foundation models alter robustness under those shifts. Models trained on controlled, in‑distribution datasets commonly suffer large performance drops when faced with seasonal, geographic, device, and sensor variability; our goal is to measure those drops and assess whether large pretrained backbones mitigate them rather than to introduce a new segmentation method.
Concretely, we analyze segmentation performance on a large, multi‑year field dataset collected in Germany and Spain (2018–2020) covering 14 plant species and 4 damage classes, and evaluate generalization across three real‑world shift axes: temporal/device drift (2023), geographic transfer to the United States, and extreme sensor/viewpoint shift to drone imagery (2024). We compare a foundation‑model backbone (DINOv2) combined with hierarchical taxonomic inference against prior baselines to isolate the impact of pretrained representations on robustness.
Our evaluation shows foundation models substantially reduce degradation from domain drift: species‑level F1 increases from 0.52 to 0.87 on in‑distribution data, and foundation backbones retain far more performance under shift (moderate shift: 0.77 vs. 0.24; extreme aerial shift: 0.44 vs. 0.14). Hierarchical taxonomic inference further cushions failures by producing useful coarser‑level predictions when species‑level labels break down (e.g., family F1: 0.68, class F1: 0.88 on aerial imagery). Error analysis reveals that failures under severe shift stem primarily from vegetation–soil confusion, suggesting that taxonomic distinctions remain preserved despite background and viewpoint variability.
These results provide empirical evidence that foundation visual models meaningfully mitigate domain drift in operational phenotyping pipelines, and they offer practical guidance for deploying robust segmentation systems in herbicide research trials across diverse regions and sensing setups.
@@ 7 Conclusions
This study evaluates the effect of domain drift on multi‑species plant segmentation and herbicide‑damage identification, and quantifies how self‑supervised vision foundation models (DINOv2) change robustness in realistic agricultural scenarios. Rather than proposing a novel segmentation algorithm, our emphasis is empirical: we measure performance degradation across seasons, geographies, devices, and sensor modalities, and test whether large pretrained visual representations reduce those degradations in operational phenotyping contexts. Experiments span multiple years and platforms (smartphones, handheld cameras, and drones) and include herbicide field trials covering more than 70 species, exposing models to substantial variability in acquisition conditions.
The three‑stage evaluation delivers three core findings. First, models using a DINOv2 backbone substantially outperform a DeepLabV3+ baseline on in‑distribution data (species F1: 0.87 vs. 0.52) and retain more performance under realistic shifts (moderate temporal/device shift: 0.77 vs. 0.24; extreme aerial shift: 0.44 vs. 0.14). Second, hierarchical taxonomic inference provides useful robustness: when fine‑grained species predictions degrade (particularly on aerial imagery), coarser taxonomic outputs (family F1: 0.68, class F1: 0.88) remain informative for downstream phenotyping tasks. Third, remaining failures under the most severe shifts are dominated by vegetation–soil separation and viewpoint changes rather than wholesale loss of species‑specific cues, suggesting pretrained features preserve relevant biological distinctions even as background and scale vary.
Deployment within BASF’s phenotyping pipeline offers practical confirmation that foundation‑backboned models can improve operational performance compared to conventional baselines. That said, the training distribution remains concentrated in Germany and Spain (2018–2020), which likely contributes to reduced absolute performance at U.S. sites despite consistent relative gains versus the baseline.
Limitations point to clear, empirical next steps. Damage classification for early or rare symptom patterns remains challenging (e.g., initial damage F1: 0.28; uncommon damage types F1: 0.05), and drone imagery shows persistent vegetation–soil confusion; both issues likely reflect annotation coverage and representational gaps for aerial viewpoints. Incorporating more aerial and geographically diverse training data, standardizing damage annotations, and adding uncertainty estimation would directly address observed failure modes and improve automated quality control.
In summary, our systematic cross‑domain evaluation demonstrates that self‑supervised foundation visual models materially mitigate the effects of domain drift in complex agricultural settings. The quantitative results and error analyses here are intended as a practical guide for deploying robust segmentation systems in herbicide research trials: they show where pretrained representations help most, where failures persist, and which dataset expansions or controls are likely to yield the greatest operational benefit.